Solve the differential equation by variation of parameters.

y'' + 3y' + 2y = 1/ 1 + e^x

y(x) =____
1
Expert's answer
2020-07-01T17:29:21-0400

y'' + 3y' + 2y = 1/( 1 + ex )

Solution:

y'' + 3y' + 2y =0

k2+3k+2=0

k1=-1,k2=-2

y0(x)=c1e-x+c2e-2x

y'' + 3y' + 2y = 1/( 1 + ex )

y(x)=c1(x)e-x+c2(x)e-2x

Let

c1'(x)e-x+c2'(x)e-2x=0,

-c1'(x)e-x-2c2'(x)e-2x=1/( 1 + ex).

Then

c2'(x)=-c1'(x)ex => c1'(x)=ex/(1+ex) => c1(x)="\int"ex/(1+ex) dx="\int" d(1+ex)/(1+exl)=ln(1+ex)+c1

c2'(x)=-e2x/(1+ex)=> c2(x)= -∫e2x/(1+ex) dx=-"\int"(ex+1-1)d(1+ex)/(1+ex)=-"\int" (1-1/(1+ex))d(1+ex)=

=-(1+ex)+ln(1+ex)+c2

y(x)=c1(x)e-x+c2(x)e-2x=(ln(1+ex)+c1)e-x+(-(1+ex)+ln(1+ex)+c2)e-2x=

=(c1-1)e-x+(c2-1)e-2x+e-xln(1+ex)+e-2xln(1+ex).

Let c1-1->c1, c2-1->c2.

Answer:

y(x)=c1e-x+c2e-2x+e-xln(1+ex)+e-2xln(1+ex).